# Using Dynamic Safety Margins as Control Barrier Functions

### v2

@@ 3 From Dynamic Safety Margins to Control Barrier Functions
Based on [33], we can formulate a reduced set of sufficient conditions that guarantee $\beta$ is locally Lipschitz continuous.
Theorem 3.18.
[33][Theorem 1] Consider $\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ defined as in (26) and suppose the following conditions hold at a point $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})\in\mathcal{C}$:
  1. 1.
    $\omega^{*}(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})>0$,
  2. 2.
    $\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}$ and $\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}$ are locally Lipschitz continuous at $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})$.
Then, the feedback $\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is unique and locally Lipschitz continuous at $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})$.
The following lemma guarantees that $\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$ wherever the feasible set has a nonempty interior.
@@
If $\operatorname{Int}\left(\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)\neq\emptyset$, then $\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$.
Proof 3.20.
Note that $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathbb{U}\times\mathbb{R}^{l}:A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})[\boldsymbol{\mathbf{u}};~{}\boldsymbol{\mathbf{w}}]\leq b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\}$. Let $(\boldsymbol{\mathbf{u}}_{0},\boldsymbol{\mathbf{w}}_{0})\in\operatorname{Int}\left(\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)$, that is, $A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})[\boldsymbol{\mathbf{u}}_{0};\boldsymbol{\mathbf{w}}_{0}]<b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. It follows that $\exists\omega_{0}>0$ such that $A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})[\boldsymbol{\mathbf{u}}_{0};\boldsymbol{\mathbf{w}}_{0}]+\boldsymbol{\mathbf{1}}\omega_{0}\leq b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. The point $(\boldsymbol{\mathbf{u}}_{0},\boldsymbol{\mathbf{w}}_{0},\omega_{0})$ is feasible in the LP defining $\omega^{*}$. Thus, $\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\omega_{0}>0$.
4 Control-sharing Property
Often, the sets $\mathcal{X}$ and $\mathcal{U}$ are the intersection of multiple state and input constraints. In these cases, it is easier to design a separate CBF for each constraint and then enforce all the conditions in one CBF-based program. Unfortunately, there is typically no guarantee that feasibility is retained when all constraints are superimposed. This section shows that DSMs designed for the same prestabilized system have the control-sharing property [9], which makes their CBF conditions compatible.
Consider again the control-affine system (1). Given $q\in\mathbb{N}$ distinct CBFs $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$ and their associated control invariant sets $\mathcal{C}_{i}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{i}\mid h_{i}(\boldsymbol{\mathbf{x}})\geq 0\}$, define the $i$-th set of safe inputs as
$\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})=\left\{\boldsymbol{\mathbf{u}}\in\mathcal{U}\mid L_{f}h_{i}(\boldsymbol{\mathbf{x}})+L_{g}h_{i}(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}+\alpha_{i}\big{(}h_{i}(\boldsymbol{\mathbf{x}})\big{)}\geq 0\right\},$
where each $\alpha_{i}$ is a class $\mathscr{K}_{\infty}$ function associated with $h_{i}$. If the conditions of Lemma 1 hold for each $h_{i}$, we have that
$\forall i\in\{1,\ldots,q\},\quad\boldsymbol{\mathbf{x}}\in\mathcal{C}_{i}\implies\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})\neq\emptyset.$ (28)
However, there is generally no guarantee that the intersection $\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})$ is nonempty.
Definition 4.21.
[9, Definition 3] Consider $q\in\mathbb{N}$ CBFs $h_{i}:\mathcal{D}_{i}\to\mathbb{R}$. These CBFs are said to have the control-sharing property if
$\forall\boldsymbol{\mathbf{x}}\in\bigcap_{i=1}^{q}\mathcal{C}_{i},\quad\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}})\neq\emptyset.$ (29)
If the CBFs $h_{i}$ have the control-sharing property, then the optimization-based control policy $\beta$ defined in (12) with
$\mathcal{C}=\bigcap_{i=1}^{q}\mathcal{C}_{i},\quad\mathcal{K}(\boldsymbol{\mathbf{x}})=\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}}),$ (30)
is feasible for all $\boldsymbol{\mathbf{x}}\in\mathcal{C}$. If it is also locally Lipschitz continuous, then $\boldsymbol{\mathbf{x}}(t)\in\mathcal{C}$ for all $t\geq 0$.
Now, suppose the state $\mathcal{X}$ and input $\mathcal{U}$ constraint sets can be decomposed into $q\in\mathbb{N}$ closed constraint sets $\mathcal{X}_{i}\subset\mathbb{X}$ and $\mathcal{U}_{i}\subset\mathbb{U}$ such that $\mathcal{X}\times\mathcal{U}\subset\mathcal{X}_{i}\times\mathcal{U}_{i},~{}\forall i\in\{1,\ldots,q\},$ and $\bigcap_{i=1}^{q}\mathcal{X}_{i}\times\mathcal{U}_{i}=\mathcal{X}\times\mathcal{U}$. The $i$-th steady-state admissible set is $\mathcal{V}_{i}=\{\boldsymbol{\mathbf{v}}\in\mathbb{V}\mid\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{i},~{}\bar{u}(\boldsymbol{\mathbf{r}})\in\mathcal{U}_{i}\}$ and note that $\bigcap_{i=1}^{q}\mathcal{V}_{i}=\mathcal{V}$. Let $\pi:\mathbb{X}\times\mathbb{V}\to\mathbb{U}$ be a prestabilizing controller as described above and assume that $\Delta_{i}:\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\to\mathbb{R}$ are dynamic safety margins for $\pi$ with respect to the constraint sets $\mathcal{X}_{i}$, $\mathcal{U}_{i}$. The following theorem shows that these DSMs have the control-sharing property.
Theorem 4.22.
Any number of DSMs $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for the same prestabilizing policy $\pi$ have the control-sharing property.
Proof 4.23.
Let $q\in\mathbb{N}$ and consider $q$ DSMs, $\Delta_{i}$, for $\pi$. If $\mathcal{C}=\bigcap_{i=1}^{q}\mathcal{C}_{i}=\emptyset$, Definition 4.21 holds trivially. Otherwise, let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{C}$ and $i\in\{1,\ldots,q\}$ be given. Since $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, it follows by (19b) that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}_{i}$. For any $\alpha\in\mathscr{K}_{\infty}$, it can be shown that the augmented input $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Since $i$ was arbitrary, it follows that $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\bigcap_{i=1}^{q}\mathcal{K}_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$ and we conclude that the DSMs $\Delta_{i}$ have the control-sharing property.
The next section provides a specific method for constructing DSMs using Lyapunov functions.
5 Lyapunov-based Dynamic Safety Margins
This section shows that the Lyapunov-based DSMs in [22, 28, 29, 30] can be used within the context of this paper. Let a continuously differentiable function $V:\mathbb{X}\times\mathbb{V}\to\mathbb{R}$ be a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$. That is, for all $\boldsymbol{\mathbf{v}}\in\mathbb{V}$,
$V\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}$ $=0,$ (31a)
$V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ $>0,\quad\forall\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\},$ (31b)
$\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ $\leq 0,\quad\forall\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}.$ (31c)
Additionally, we require $V$ to be strictly monotonically increasing over the sets $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Formally, we have the following assumption.
Assumption 4
Given any $\boldsymbol{\mathbf{v}}\in\mathbb{V}$, $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}}):\mathbb{X}\to\mathbb{R}$.
Recalling the set $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ defined in (18), let $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}=\mathbb{X}\setminus\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ be its complement, which is open in $\mathbb{X}$. We define $\Gamma^{*}:\mathbb{V}\to\mathbb{R}$ as
When positive, $\Gamma^{*}(\boldsymbol{\mathbf{v}})$ represents the largest constraint-admissible level set of the Lyapunov function $V$ at reference $\boldsymbol{\mathbf{v}}$. Our definition for $\Gamma^{*}$ is an extension of the threshold value in [22] to the entire reference space $\mathbb{V}$. With this, $\Gamma^{*}$ has the following property.
Lemma 5.24.
@@ 5 Lyapunov-based Dynamic Safety Margins
Let $V:\mathbb{X}\times\mathbb{V}\to\mathbb{R}$ be a reference-dependent Lyapunov function that satisfies Assumption 4. We have that
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0\quad\iff\quad\boldsymbol{\mathbf{v}}\in\mathcal{V}.$ (33)
Proof 5.25.
@@
For the other direction, we show the contrapositive statement. Let $\boldsymbol{\mathbf{v}}\in\mathbb{V}\setminus\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\notin\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, and $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Since $\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{X}$, there exists $\delta>0$ such that $\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}=\{\boldsymbol{\mathbf{x}}\in\mathbb{X}:\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|<\delta\}\subset\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Furthermore, by compactness of $\partial\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, there exists $\boldsymbol{\mathbf{x}}_{0}\in\partial\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, with $\boldsymbol{\mathbf{x}}_{0}\neq\bar{x}(\boldsymbol{\mathbf{v}})$ because $\bar{x}(\boldsymbol{\mathbf{v}})\notin\partial\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, such that
$\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{B}(\bar{x}(\boldsymbol{\mathbf{v}}),\delta)}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})>0.$ (34)
By Assumption 4, $V$ is monotonically increasing over $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, and its level sets contained in $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ are connected. It follows that $\mathcal{L}\triangleq\{\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}:V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})\}\subset\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$. Now, let $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ be given and note that $\boldsymbol{\mathbf{x}}\notin\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}$, implying that $\boldsymbol{\mathbf{x}}\notin\mathcal{L}$. So, we must have that $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})$. As a result,
$\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})>0,$ (35)
and we conclude that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq-V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}})<0$.
Let us define the stability threshold value
@@
Let $V:\mathbb{X}\times\mathbb{V}\to\mathbb{R}$ be a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ satisfying Assumption 4. Let $\Gamma:\mathbb{V}\to\mathbb{R}$ be a continuously differentiable function such that $\forall\boldsymbol{\mathbf{v}}\in\mathcal{V}$, $0\leq\Gamma(\boldsymbol{\mathbf{v}})\leq\min\left(\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\right)$ and $\forall\boldsymbol{\mathbf{v}}\in\mathbb{V}\setminus\mathcal{V}$, $\Gamma(\boldsymbol{\mathbf{v}})<0$. Then, $\Delta:\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\to\mathbb{R}$, defined as
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\Gamma(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (37)
@@
Proof 5.27.
To prove property (19a) in Definition 2.4, let $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Then, by Lemma 5.24, $\Delta\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\Gamma(\boldsymbol{\mathbf{v}})-V\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big{)}=\Gamma(\boldsymbol{\mathbf{v}})\geq 0$.
@@
To prove property (19b), let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ be such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. For a contradiction, assume $\boldsymbol{\mathbf{x}}\notin\mathcal{X}_{\boldsymbol{\mathbf{v}}}$. Then
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$
Additionally, $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$ implies that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq\Gamma(\boldsymbol{\mathbf{v}})\geq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, which further implies by Lemma 5.24 that $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Thus,
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\Gamma(\boldsymbol{\mathbf{v}})\leq\Gamma^{*}(\boldsymbol{\mathbf{v}}).$ (38)

### v3

@@ 3.3 Existence and Uniqueness of Solutions
exists and is unique. Based on [38], we can formulate a reduced set of sufficient conditions that guarantee $\beta$ is locally Lipschitz continuous.
Theorem 3.19 ([38, Theorem 5.2]).
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$, let $\rho$ be a navigation field and let $\eta>0$. Assume the input constraints are polyhedral $\mathcal{U}=\{\boldsymbol{\mathbf{u}}\in\mathbb{R}^{m}:M\boldsymbol{\mathbf{u}}\leq\boldsymbol{\mathbf{b}}\}$, and the functions $\pi(\cdot,\boldsymbol{\mathbf{r}})$ and $\rho(\cdot,\boldsymbol{\mathbf{r}})$ are continuously differentiable. If these constraint qualifications (CQs) hold
  1. a)
    (Slater CQ) There exists $\alpha\in\mathscr{K}_{\infty}$ such that for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\operatorname{Int}\left(\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)\neq\emptyset$;
  2. b)
    (Constant Rank CQ) There exists a neighborhood of every point in $\tilde{\mathcal{C}}$ wherein, for any subset $\mathcal{I}\subset\{1,\ldots,n_{c}\}$, the number of linearly independent vectors in $\left\{\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix},i\in\mathcal{I}\right\}$ remains constant;
then the optimization-based policy $\beta$ is locally Lipschitz continuous and the solution to the closed-loop system (32) exists and is unique.
Remark 3.20.
In the special case that $n_{c}=1$, both Slater and Constant Rank CQs hold if $\Delta$ has uniform relative degree $1$ over $\tilde{\mathcal{C}}$ and $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\operatorname{Int}\left(\mathcal{U}\right)$ for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$.
4 Lyapunov-based Dynamic Safety Margins
This section shows that the Lyapunov-based DSMs in [23, 30, 31, 32] can be used within the context of this paper.
Definition 4.21.
A continuously differentiable function $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ is a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ if there exist functions $\gamma_{1},\gamma_{2}\in\mathscr{K}_{\infty}$ and $\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, there exists a neighborhood of $\bar{x}(\boldsymbol{\mathbf{v}})$, denoted $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$, wherein
$\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)}$ $\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\gamma_{2}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)},$ (33a)
$\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ $\leq 0.$ (33b)
Remark 4.22.
Note that the reference-dependent Lyapunov function $V$ needs not certify asymptotic stability of the equilibria of the prestabilized system $f_{\pi}$. That is, $\dot{V}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$ is acceptable for $\boldsymbol{\mathbf{x}}\neq\bar{x}(\boldsymbol{\mathbf{v}})$. This allows for more general classes of Lyapunov functions to be considered and ultimately simplifies the design of the DSM.
Additionally, we require $V$ has non-vanishing gradient near the equilibrium points. Formally, we have the following assumption.
Assumption 4
Given any $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}}):\mathbb{R}^{n}\to\mathbb{R}$.
Recalling the reference-dependent state constraint set $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ defined in (19), let $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}=\mathbb{R}^{n}\setminus\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ be its complement, which is open in $\mathbb{R}^{n}$. We define the safety threshold value $\Gamma^{*}:\mathbb{R}^{l}\to\mathbb{R}$
When positive, $\Gamma^{*}(\boldsymbol{\mathbf{v}})$ represents the largest constraint-admissible level set of the Lyapunov function $V$ at reference $\boldsymbol{\mathbf{v}}$. Our definition for $\Gamma^{*}$ is an extension of the threshold value in [23] to the entire reference space $\mathbb{R}^{l}$. With this, $\Gamma^{*}$ has the following property.
@@ 4 Lyapunov-based Dynamic Safety Margins
Let $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ be a reference-dependent Lyapunov function that satisfies Assumption 4. We have that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0$ if and only if $\boldsymbol{\mathbf{v}}\in\mathcal{V}$.
Proof 4.24.
@@
For the other direction, we show the contrapositive statement. Let $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}\setminus\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Since $\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{R}^{n}$, there exists $\delta>0$ such that $\mathcal{B}\big{(}\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big{)}=\{\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}\mid\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|<\delta\}\subset\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Thus,
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq-\gamma_{1}(\delta)<0.$ (35)
Let us also define the stability threshold value $\overline{\Gamma}:\mathbb{R}^{l}\to\mathbb{R}$
@@
Let $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ be a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ satisfying Assumption 4. If $\Gamma^{*}:\mathbb{R}^{l}\to\mathbb{R}$ and $\overline{\Gamma}:\mathbb{R}^{l}\to\mathbb{R}$ are continuously differentiable, and $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{V}\subset\mathbb{R}^{l}$ are compact, then for any $\epsilon\in(0,1)$,
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\begin{bmatrix}\Gamma^{*}(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ (1-\epsilon)\overline{\Gamma}(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\end{bmatrix},$ (37)
@@
Proof 4.26.
@@
To prove property (22a), let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{D}}$. Then, $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, which implies that $\boldsymbol{\mathbf{v}}\in\mathcal{V}$ by Lemma 4.23. Next, for a contradiction, assume $\boldsymbol{\mathbf{x}}\notin\mathcal{X}_{\boldsymbol{\mathbf{v}}}$. Then
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$